### v1

# Sustainable Exploitation Equilibria for Dynamic Games

Abstract
We introduce the Sustainable Exploitation Equilibrium (SEE), a refinement of Markov Perfect Equilibrium (MPE) for dynamic games with an exploiter–exploitee structure. SEE imposes two additional discipline conditions: (i) viability, requiring state trajectories to remain inside a sustainability set; and (ii) renegotiation-proofness with exploiter-optimal selection, to retain only those viable equilibria that are immune to Pareto-improving renegotiations, with ties resolved in favor of the exploiter. In our base formulation the exploitee cannot exit the relationship (no outside option), but retains a strategic effort margin that affects dynamics and payoffs. We establish existence under appropriate conditions and illustrate SEE in a hegemon-client model of foreign politics, where tribute demands trade off against the client’s governance effort.
Keywords: Markov perfect equilibrium, equilibrium refinements, viability, renegotiation-proofness, sustainability
1 Introduction
Exploitation arises when one agent extracts value from another whose continued functioning is essential for the relationship itself. Standard equilibrium concepts such as Markov Perfect Equilibrium (MPE) may admit outcomes in which the exploiter extracts too aggressively, driving the state outside the range in which the exploitee can survive. To address this, the paper introduces the Sustainable Exploitation Equilibrium (SEE), a refinement of stationary MPE that selects only those equilibria that (i) keep the state within a viability set and (ii) are immune to Pareto-improving renegotiations, with exploiter-optimal selection. The contribution is to formalize a simple and tractable equilibrium notion suited to persistent asymmetric relationships in which collapse is feasible but avoided in equilibrium. SEE is motivated by settings where one party cannot exit but remains responsive, such as in hegemon–client dynamics, as the exploitee retains a strategic effort margin that shapes both payoffs and dynamics. In this way, SEE formalizes the idea of “sustainable exploitation”: the exploiter extracts as much as possible while ensuring the long-run survival and productivity of the exploitee.
2 Related Literature
Our work connects three strands of existing literature.
Equilibrium refinements.
SEE builds on the refinement tradition in dynamic games. (MaskinTirole2001a) develop the theory of Markov perfect equilibrium; (Bernheim1987) introduce coalition-proof Nash equilibrium; (Selten1975) trembling-hand perfection; and (FarrellMaskin1989) renegotiation-proofness in repeated games.
Viability and dynamic control.
The viability constraint draws from (Aubin1991) Viability Theory and constrained Markov decision processes (Altman1999), ensuring state trajectories remain within sustainability sets. This connects to the resource economics literature, e.g. (Dasgupta1979), (Chichilnisky1997), and (Dockner2000), where long-run viability is central to dynamic exploitation of resources.
Political economy and relational contracts.
SEE also relates to political economy models of extraction and state capacity ((AcemogluRobinson2006); (BesleyPersson2011)) and to relational contracting without full commitment ((ThomasWorrall1988); (Levin2003)). In these contexts, weaker parties cannot exit but retain effort margins that shape dynamics. SEE formalizes the logic of “restrained exploitation” within a Markovian framework.
3 Model and Definition
Environment and timing.
Time is discrete, $t=0,1,2,\dots$. The state $s_{t}\in S$ evolves on a compact metric space $S$. In each period:
  1. 1.
    The exploiter $X$ observes $s_{t}$ and chooses extraction $x_{t}\in X(s_{t})$.
  2. 2.
    The exploitee $E$ observes $(s_{t},x_{t})$ and chooses effort $e_{t}\in E(s_{t})$.
  3. 3.
    The state transitions to $s_{t+1}\sim Q(\cdot\mid s_{t},x_{t},e_{t})$ (deterministic dynamics are $s_{t+1}=f(s_{t},x_{t},e_{t})$).
Per-period payoffs are $u^{X}(s_{t},x_{t},e_{t})$ and $u^{E}(s_{t},x_{t},e_{t})$, bounded and continuous. Both players discount with $\delta\in(0,1)$.
Roles of the players
Definition 1 (Exploiter).
The exploiter $X$ directly extracts value via $x\in X(s)$, appropriating resources and typically reducing the future state. $X$’s stage payoff $u^{X}(s,x,e)$ is weakly increasing in $x$ for fixed $(s,e)$. A Markov policy for $X$ is $\sigma^{X}:S\to X(s)$.
Definition 2 (Exploitee (no exit, effort choice)).
The exploitee $E$ cannot exit the relationship (no outside option) but chooses effort $e\in E(s)$ after observing $(s,x)$. Effort affects contemporaneous payoffs and/or the law of motion. A Markov policy for $E$ is $\sigma^{E}:S\times X(s)\to E(s)$ (second-mover response).
Equilibrium notion (Markov–Stackelberg Equilibrium).
A Markov–Stackelberg Equilibrium (MSE) consists of policies $(\sigma^{X},\sigma^{E})$ such that: (i) for each $(s,x)$, $\sigma^{E}(s,x)$ is a best response to $\sigma^{X}$ given follower timing; (ii) $\sigma^{X}(s)$ anticipates $\sigma^{E}$ and is a best response given $E$’s continuation behavior. Values $(W^{X},W^{E})$ solve the associated Bellman system.
Definition 3 (Sustainable Exploitation Equilibrium (SEE)).
Let $V\subseteq S$ be a nonempty compact viability set that is invariant under admissible state transitions. Let $\mathcal{E}$ denote the set of all stationary Markov Perfect Equilibria, and let $\mathcal{E}_{V}\subseteq\mathcal{E}$ denote the subset whose induced state process remains in $V$. We assume $\mathcal{E}_{V}\neq\emptyset$.
A Sustainable Exploitation Equilibrium is any stationary MPE $\sigma^{*}\in\mathcal{E}$ satisfying:
  1. (i)
    Viability: The state induced by $\sigma^{*}$ remains in the viability set $V$ for all $s\in V$ (i.e., $\sigma^{*}\in\mathcal{E}_{V}$).
  2. (ii)
    Renegotiation-proofness: At any state $s\in V$, there is no alternative viable profile $\tilde{\sigma}\in\mathcal{E}_{V}$ such that:
    $W_{E}(\tilde{\sigma};s)\geq W_{E}(\sigma^{*};s)\quad\text{and}\quad W_{X}(\tilde{\sigma};s)\geq W_{X}(\sigma^{*};s),$
    with at least one inequality strict.
  3. (iii)
    Exploiter-optimality: Among all renegotiation-proof elements of $\mathcal{E}_{V}$, $\sigma^{*}$ maximizes the exploiter’s value $W_{X}(\cdot;s)$ at the state $s$ (with ties broken lexicographically).
Any profile satisfying (i)–(iii) is a Sustainable Exploitation Equilibrium.
4 Relation to Nash and Markov Perfect Equilibrium
We position the Sustainable Exploitation Equilibrium (SEE) within the standard equilibrium hierarchy.
Nash equilibrium (NE).
A strategy profile $\sigma=(\sigma_{X},\sigma_{E})$ is a Nash equilibrium if no player can profitably deviate unilaterally. NE is broad and may rely on history-dependent strategies.
Subgame perfect equilibrium (SPE).
A refinement of NE, SPE requires sequential rationality in every subgame. This eliminates non-credible threats. Clearly,
$\text{SPE}\subseteq\text{NE}.$
Markov perfect equilibrium (MPE).
A further refinement of SPE for dynamic games with a payoff-relevant state $s\in S$. In MPE, strategies are Markovian: $\sigma_{i}(s)$ depends only on the current state, not the full history. For two-player sequential-move environments, the relevant benchmark is a Markov–Stackelberg equilibrium (MSE): the follower chooses $\sigma_{E}(s,x)$ as a best response given the state and leader’s action, while the leader anticipates this response in selecting $\sigma_{X}(s)$. Thus,
$\text{MPE (MSE)}\subseteq\text{SPE}\subseteq\text{NE}.$
Sustainable Exploitation Equilibrium (SEE).
SEE is defined as a viable, renegotiation-proof MPE with exploiter-optimal selection. Formally, a stationary Markov profile $(\sigma_{X},\sigma_{E})$ is an SEE if the aforementioned conditions of Definition 3 are satisfied.
It follows immediately that
$\text{SEE}\subseteq\text{MPE}\subseteq\text{SPE}\subseteq\text{NE}.$
5 Existence
Theorem 1 (Existence of SEE).
Suppose the state space $S$ and action sets $X(s),E(s)$ are compact; payoffs are bounded and continuous; and the transition kernel $Q$ satisfies the Feller property. Let $V\subseteq S$ be a nonempty compact viability set that is invariant under admissible state transitions. Assume that there exists at least one stationary Markov Perfect Equilibrium whose induced state process remains in $V$. Then:
(i) the set of viable stationary Markov Perfect Equilibria is nonempty and compact;
(ii) the subset of renegotiation-proof viable equilibria is nonempty; and
(iii) an exploiter-optimal selection exists.
Hence a Sustainable Exploitation Equilibrium (SEE) exists.
Sketch.
(1) Existence of stationary MPE. Under compact state and action spaces, bounded and continuous payoffs, and Feller continuity of the transition kernel, the stationary best-response correspondence is nonempty, convex-valued, and upper hemicontinuous. By Glicksberg’s fixed point theorem, there exists at least one stationary Markov Perfect Equilibrium (Maskin and Tirole, 2001).
(2) Viability. Let $\mathcal{E}$ denote the set of stationary MPE and define $\mathcal{E}_{V}\subseteq\mathcal{E}$ as the subset whose induced state process remains in the invariant compact set $V$. By assumption, $\mathcal{E}_{V}$ is nonempty. Compactness of $S$ and continuity of payoffs and transitions imply that the associated set of equilibrium value functions for $\mathcal{E}_{V}$ is compact in the product topology.
(3) Renegotiation-proofness. The set of feasible continuation values generated by $\mathcal{E}_{V}$ is compact. Therefore there exists at least one equilibrium in $\mathcal{E}_{V}$ that is not strictly Pareto-dominated by any other viable equilibrium; such an equilibrium is renegotiation-proof.
(4) Selection. The exploiter’s value function is continuous on the nonempty compact set of renegotiation-proof viable equilibria and hence attains a maximum. Any maximizer is a Sustainable Exploitation Equilibrium. ∎
6 Example: Foreign Politics (Hegemon–Client SEE)
6.1 Primitives and timing
Two players: a hegemon $H$ (the exploiter) and a client state $C$ (the exploitee). The state $s_{t}\in S\subset\mathbb{R}_{+}$ represents client political capacity/stability (e.g., tax capacity, administrative control, consent). Time is discrete $t=0,1,2,\dots$.
Within each period:
  1. 1.
    $H$ observes $s_{t}$ and chooses an extraction/pressure level $x_{t}\in X(s_{t})$ (e.g., tribute, resource concessions, policy compliance).
  2. 2.
    $C$ observes $(s_{t},x_{t})$ and chooses effort $e_{t}\in E(s_{t})$ (e.g., governance effort, compliance/implementation, regime maintenance).
  3. 3.
    The state updates to $s_{t+1}=f(s_{t},e_{t})-h(x_{t})$, where $f$ is increasing in $s$ and $e$ and $h$ is increasing in $x$; or more generally $s_{t+1}\sim Q(\cdot\mid s_{t},x_{t},e_{t})$.
Both discount with factor $\delta\in(0,1)$. The client cannot exit the relationship (no outside option) but has strategic effort. A viability requirement $s_{t}\geq s_{\min}$ avoids regime collapse or state failure.
6.2 Payoffs and feasibility
Per-period payoffs are continuous and bounded:
$u^{H}(s,x,e)\;=\;\pi(x)\;-\;k(s,e),\qquad u^{C}(s,x,e)\;=\;b(s)\;-\;\phi(e)\;-\;d(x),$
with $\pi^{\prime}(x)>0$, $k\geq 0$ (costs to monitor/coerce or instability spillovers), $b^{\prime}(s)\geq 0$, $\phi$ convex with $\phi^{\prime}(0)=0$, and $d^{\prime}(x)\geq 0$. Feasibility (viability) imposes
$x\;\leq\;f(s,e)-s_{\min}.$
6.3 Equilibrium notion and SEE
We adopt a Markov–Stackelberg Equilibrium (MSE): the follower $C$ best responds with $\sigma^{C}(s,x)$, and the leader $H$ best responds with $\sigma^{H}(s)$ anticipating $\sigma^{C}$. A Sustainable Exploitation Equilibrium (SEE) is a viable, renegotiation-proof MSE with exploiter-optimal selection (as in Definition 1).
6.4 Follower’s best response
Given $(s,x)$ and anticipating future play, $C$ solves
$\max_{e\in E(s)}\;\Big\{b(s)-\phi(e)-d(x)+\delta W^{C}\!\big(f(s,e)-h(x)\big)\Big\}.$
An interior first-order condition is
$\phi^{\prime}(e^{*})\;+\;\delta\,W^{C}_{s}\!\big(s^{\prime}\big)\,f_{e}(s,e^{*})\;=\;0,\qquad s^{\prime}\equiv f(s,e^{*})-h(x).$
Under standard concavity/regularity, the implicit function theorem yields $\partial e^{*}/\partial x$.
6.5 Leader’s SEE problem
Anticipating $e^{*}(s,x)$, the hegemon solves
$V^{H}(s)\;=\;\max_{x\in X(s)}\;\Big\{\pi(x)-k(s,e^{*}(s,x))+\delta\,V^{H}\!\big(f(s,e^{*}(s,x))-h(x)\big)\Big\}$
subject to viability $x\leq f(s,e^{*}(s,x))-s_{\min}$. Let $\mu\geq 0$ be the multiplier on viability. The Lagrangian is
$\mathcal{L}(x;s,\mu)\;=\;\pi(x)-k(s,e^{*})+\delta V^{H}(s^{\prime})\;+\;\mu\big(f(s,e^{*})-h(x)-s_{\min}\big),$
where $e^{*}\equiv e^{*}(s,x)$ and $s^{\prime}\equiv f(s,e^{*})-h(x)$.
The (interior) first-order condition with envelope terms from $e^{*}(s,x)$ is
$\pi^{\prime}(x)\;-\;\delta V^{H}_{s}(s^{\prime})\,\big(h^{\prime}(x)-f_{e}(s,e^{*})\,\partial e^{*}/\partial x\big)\;-\;\mu\,\big(h^{\prime}(x)-f_{e}(s,e^{*})\,\partial e^{*}/\partial x\big)\;=\;0,$
with complementary slackness $\mu\big(f(s,e^{*})-h(x)-s_{\min}\big)=0$.
Interpretation.
The hegemon equates marginal extraction benefits $\pi^{\prime}(x)$ to the discounted shadow cost of lost future capacity $h^{\prime}(x)$ net of the induced effort response $f_{e}\partial e^{*}/\partial x$. When viability binds, the shadow price $\mu>0$ further restrains extraction.
6.6 Stationary SEE and regimes
A stationary SEE consists of $(x^{*}(s),e^{*}(s,x))$ and an invariant distribution supported on $V=\{s\geq s_{\min}\}$. Two regimes arise:
  • Interior (viability slack, $\mu=0$). If the unconstrained fixed point $s^{*}$ of $s^{\prime}=f(s,e^{*}(s,x^{*}))-h(x^{*})$ satisfies $s^{*}>s_{\min}$, the FOCs imply a generalized “golden rule” where effective regeneration $f_{s}+f_{e}\,\partial e^{*}/\partial s$ balances discounting (details depend on primitives).
  • Boundary (viability binds, $\mu>0$). If the unconstrained $s^{*}$ would fall below $s_{\min}$, SEE selects $s^{*}=s_{\min}$ and $x^{*}=f(s_{\min},e^{*}(s_{\min},x^{*}))-s_{\min}$, with $\mu$ set to satisfy the FOC.
Additional Remark on Outside Option Expansions
One might alternatively assume that the exploitee $E$ has an outside option $U^{out}_{E}(s)$, in which case viability must be supplemented by an individual rationality (IR) constraint: $W^{E}(s)\geq U^{out}_{E}(s)$ for all $s\in V$. This yields an even tighter refinement, as it restricts feasible equilibria to those in which the exploitee weakly prefers participation over exit. We deliberately exclude outside options in the baseline SEE definition, since the core notion of exploitation is precisely that the exploitee cannot walk away. In other settings, the SEE framework can be adapted straightforwardly by adding the IR condition, producing a refinement that nests our baseline definition, but applies to a broader class of sustainable strategic relationships.
7 Conclusion
With a non-exiting but strategically responsive exploitee, SEE refines MPE by imposing viability and renegotiation-proof exploiter-optimal selection. In the hegemon–client example, SEE yields restrained extraction that keeps the regime just at the viability boundary. More generally, the framework applies whenever one party cannot exit but retains an effort margin, including colonial settings, labor relationships and locked-in suppliers. By combining dynamic sustainability with renegotiation-proofness, SEE offers a compact refinement of MPE for analyzing persistent exploitative interactions in economics and political economy.

### v2

@@
 We introduce the Sustainable Exploitation Equilibrium (SEE), a refinement of Markov Perfect Equilibrium (MPE) for dynamic games with an exploiter–exploitee structure.
-SEE imposes two additional discipline conditions: (i) viability, requiring state trajectories to remain inside a sustainability set; and (ii) renegotiation-proofness with exploiter-optimal selection, to retain only those viable equilibria that are immune to Pareto-improving renegotiations, with ties resolved in favor of the exploiter.
-In our base formulation the exploitee cannot exit the relationship (no outside option), but retains a strategic effort margin that affects dynamics and payoffs.
-We establish existence under appropriate conditions and illustrate SEE in a hegemon-client model of foreign politics, where tribute demands trade off against the client’s governance effort.
+SEE imposes renegotiation-proof exploiter-optimal selection on the set of rationally viable stationary Markov equilibria, where viability follows from sequential rationality when exiting a sustainability set entails catastrophic losses.
+Unlike MPE, SEE rules out equilibria in which the exploiter optimally drives the state to collapse despite positive continuation payoffs.
+The exploitee cannot exit, but retains a strategic effort margin affecting dynamics and payoffs.
+We establish existence under standard conditions, and the refinement is illustrated in a hegemon–client model of foreign politics.
 Keywords: Markov perfect equilibrium, equilibrium refinements, viability, renegotiation-proofness, sustainability
+JEL: C72, C73, D74.
 1 Introduction
 Exploitation arises when one agent extracts value from another whose continued functioning is essential for the relationship itself.
-Standard equilibrium concepts such as Markov Perfect Equilibrium (MPE) may admit outcomes in which the exploiter extracts too aggressively, driving the state outside the range in which the exploitee can survive.
-To address this, the paper introduces the Sustainable Exploitation Equilibrium (SEE), a refinement of stationary MPE that selects only those equilibria that (i) keep the state within a viability set and (ii) are immune to Pareto-improving renegotiations, with exploiter-optimal selection.
-The contribution is to formalize a simple and tractable equilibrium notion suited to persistent asymmetric relationships in which collapse is feasible but avoided in equilibrium.
+In such dynamic environments, standard equilibrium concepts such as Markov Perfect Equilibrium (MPE) may admit outcomes in which the exploiter optimally extracts too aggressively, driving the state into collapse despite positive continuation payoffs.
+These equilibria rely on the exploiter being indifferent to catastrophic breakdown once it occurs.
+To address this, this paper introduces the Sustainable Exploitation Equilibrium (SEE), a refinement of stationary Markov Perfect Equilibrium that rules out such collapse outcomes by deriving viability from rationality.
+When exiting a sustainability region entails sufficiently large losses for the exploiter—and when a feasible safe action exists that preserves the relationship—any equilibrium placing positive probability on collapse is sequentially irrational.
+Under these conditions, restricting attention to viable stationary equilibria is without loss.
+SEE then further refines the set of rationally viable stationary MPE by imposing renegotiation-proofness and exploiter-optimal selection.
 SEE is motivated by settings where one party cannot exit but remains responsive, such as in hegemon–client dynamics, as the exploitee retains a strategic effort margin that shapes both payoffs and dynamics.
 In this way, SEE formalizes the idea of “sustainable exploitation”: the exploiter extracts as much as possible while ensuring the long-run survival and productivity of the exploitee.
+This rational-viability foundation distinguishes SEE from equilibrium selection rules that impose sustainability as an external constraint.
 2 Related Literature
@@
 Values $(W^{X},W^{E})$ solve the associated Bellman system.
+Rational interpretation of viability.
+We interpret transitions outside a sustainability set $V\subset S$ as catastrophic breakdown for the exploiter, in the sense that sufficiently small reductions in extraction can strictly improve continuation payoffs by preserving the relationship.
+Under this interpretation, equilibrium behavior that places positive probability on exiting $V$ is incompatible with sequential rationality when collapse entails sufficiently large losses.
 Definition 3 (Sustainable Exploitation Equilibrium (SEE)).
@@
 Let $\mathcal{E}$ denote the set of all stationary Markov Perfect Equilibria, and let $\mathcal{E}_{V}\subseteq\mathcal{E}$ denote the subset whose induced state process remains in $V$.
-We assume $\mathcal{E}_{V}\neq\emptyset$.
+Under rational viability (formalized below), restricting attention to $\mathcal{E}_{V}$ is without loss.
 A Sustainable Exploitation Equilibrium is any stationary MPE $\sigma^{*}\in\mathcal{E}$ satisfying:
@@
 $\text{SEE}\subseteq\text{MPE}\subseteq\text{SPE}\subseteq\text{NE}.$
-5 Existence
-Theorem 1 (Existence of SEE).
-Suppose the state space $S$ and action sets $X(s),E(s)$ are compact; payoffs are bounded and continuous; and the transition kernel $Q$ satisfies the Feller property.
-Let $V\subseteq S$ be a nonempty compact viability set that is invariant under admissible state transitions.
-Assume that there exists at least one stationary Markov Perfect Equilibrium whose induced state process remains in $V$.
-Then:
-(i) the set of viable stationary Markov Perfect Equilibria is nonempty and compact;
-(ii) the subset of renegotiation-proof viable equilibria is nonempty; and
-(iii) an exploiter-optimal selection exists.
-Hence a Sustainable Exploitation Equilibrium (SEE) exists.
+5 Existence under rational viability
+We derive viability from rationality by interpreting exits from a sustainability set as catastrophic for the exploiter.
+We interpret the catastrophe penalty $M$ as a reduced-form representation of franchise loss, regime collapse, or irreversible destruction of the exploitee following breakdown of the relationship.
+Assumption (Catastrophic collapse and safe action).
+Payoffs are bounded: $|u^{X}(s,x,e)|\leq\bar{u}<\infty$.
+Let $V\subset S$ be a nonempty compact set.
+For each $s\in V$, there exists a safe action $\underline{x}(s)\in X(s)$ such that, under the exploitee’s best response,
+$Q\!\left(V\mid s,\underline{x}(s),\sigma_{E}(s,\underline{x}(s))\right)=1.$
+Theorem 1 (Existence of SEE under rational viability).
+Suppose the state space $S$ and action sets $X(s),E(s)$ are compact, payoffs are bounded and continuous, and the transition kernel $Q$ satisfies the Feller property.
+Interpret $S\setminus V$ as a collapse region generating a catastrophic loss for the exploiter.
+Then there exists $M^{\star}<\infty$ such that, for all catastrophe penalties $M\geq M^{\star}$, every stationary Markov Perfect Equilibrium of the penalized game
+$\hat{u}^{X}(s,x,e)\equiv u^{X}(s,x,e)-M\cdot\mathbf{1}\{s^{\prime}\notin V\},\qquad s^{\prime}\sim Q(\cdot\mid s,x,e),$
+is viable.
+Moreover, among viable stationary Markov Perfect Equilibria, a nonempty compact subset of renegotiation-proof equilibria exists, and an exploiter-optimal selection exists.
+Hence a Sustainable Exploitation Equilibrium exists.
 Sketch.
-(1) Existence of stationary MPE.
-Under compact state and action spaces, bounded and continuous payoffs, and Feller continuity of the transition kernel, the stationary best-response correspondence is nonempty, convex-valued, and upper hemicontinuous.
-By Glicksberg’s fixed point theorem, there exists at least one stationary Markov Perfect Equilibrium (Maskin and Tirole, 2001).
-(2) Viability.
-Let $\mathcal{E}$ denote the set of stationary MPE and define $\mathcal{E}_{V}\subseteq\mathcal{E}$ as the subset whose induced state process remains in the invariant compact set $V$.
-By assumption, $\mathcal{E}_{V}$ is nonempty.
-Compactness of $S$ and continuity of payoffs and transitions imply that the associated set of equilibrium value functions for $\mathcal{E}_{V}$ is compact in the product topology.
-(3) Renegotiation-proofness.
-The set of feasible continuation values generated by $\mathcal{E}_{V}$ is compact.
-Therefore there exists at least one equilibrium in $\mathcal{E}_{V}$ that is not strictly Pareto-dominated by any other viable equilibrium; such an equilibrium is renegotiation-proof.
-(4) Selection.
-The exploiter’s value function is continuous on the nonempty compact set of renegotiation-proof viable equilibria and hence attains a maximum.
-Any maximizer is a Sustainable Exploitation Equilibrium.
+Under compactness and continuity, stationary Markov Perfect Equilibria exist.
+Fix any $s\in V$.
+If an equilibrium action induces $Q(S\setminus V\mid s,x,e)>0$, the exploiter’s expected one-period payoff is at most $\bar{u}-Mp$ for some $p>0$.
+Under the safe action $\underline{x}(s)$, the expected penalty is zero and the payoff is at least $\bar{u}$.
+For $M$ sufficiently large, the safe action strictly dominates any action with positive exit probability.
+By continuity and compactness, a finite threshold $M^{\star}$ exists.
+Hence any equilibrium must be viable.
+Existence of renegotiation-proof and exploiter-optimal selections follows by compactness and continuity.
 ∎
@@
 6.6 Stationary SEE and regimes
-A stationary SEE consists of $(x^{*}(s),e^{*}(s,x))$ and an invariant distribution supported on $V=\{s\geq s_{\min}\}$. Two regimes arise:
+A stationary SEE consists of $(x^{*}(s),e^{*}(s,x))$ and an invariant distribution supported on $V=\{s\geq s_{\min}\}$. Under rational viability, two canonical regimes arise:
   • Interior (viability slack, $\mu=0$).
@@
 If the unconstrained $s^{*}$ would fall below $s_{\min}$, SEE selects $s^{*}=s_{\min}$ and $x^{*}=f(s_{\min},e^{*}(s_{\min},x^{*}))-s_{\min}$, with $\mu$ set to satisfy the FOC.
-Additional Remark on Outside Option Expansions
+Additional Remark on Outside Option
 One might alternatively assume that the exploitee $E$ has an outside option $U^{out}_{E}(s)$, in which case viability must be supplemented by an individual rationality (IR) constraint: $W^{E}(s)\geq U^{out}_{E}(s)$ for all $s\in V$.
 This yields an even tighter refinement, as it restricts feasible equilibria to those in which the exploitee weakly prefers participation over exit.
-We deliberately exclude outside options in the baseline SEE definition, since the core notion of exploitation is precisely that the exploitee cannot walk away.
-In other settings, the SEE framework can be adapted straightforwardly by adding the IR condition, producing a refinement that nests our baseline definition, but applies to a broader class of sustainable strategic relationships.
 7 Conclusion
-With a non-exiting but strategically responsive exploitee, SEE refines MPE by imposing viability and renegotiation-proof exploiter-optimal selection.
-In the hegemon–client example, SEE yields restrained extraction that keeps the regime just at the viability boundary.
-More generally, the framework applies whenever one party cannot exit but retains an effort margin, including colonial settings, labor relationships and locked-in suppliers.
-By combining dynamic sustainability with renegotiation-proofness, SEE offers a compact refinement of MPE for analyzing persistent exploitative interactions in economics and political economy.
+With a non-exiting but strategically responsive exploitee, SEE refines Markov Perfect Equilibrium by deriving viability from rationality and imposing renegotiation-proof exploiter-optimal selection.
+Unlike MPE, SEE eliminates equilibria in which the exploiter optimally drives the state to collapse despite positive continuation payoffs.
+In equilibrium, SEE yields either an interior “golden-rule”–type extraction path or, when unconstrained extraction would cause collapse, a boundary outcome in which the state is pinned at the viability threshold.
+The framework applies broadly to settings in which one party cannot exit but retains an effort margin—such as colonial relationships, labor arrangements, and locked-in suppliers—providing a compact refinement for analyzing persistent exploitative interactions in economics and political economy.

### v3

@@
-Sustainable Exploitation Equilibria for Dynamic Games
+Sustainable Exploitation Equilibria for Dynamic Games with Irreversible Failure
 Abstract
-We introduce the Sustainable Exploitation Equilibrium (SEE), a refinement of Markov Perfect Equilibrium (MPE) for dynamic games with an exploiter–exploitee structure.
-SEE imposes renegotiation-proof exploiter-optimal selection on the set of rationally viable stationary Markov equilibria, where viability follows from sequential rationality when exiting a sustainability set entails catastrophic losses.
-Unlike MPE, SEE rules out equilibria in which the exploiter optimally drives the state to collapse despite positive continuation payoffs.
-The exploitee cannot exit, but retains a strategic effort margin affecting dynamics and payoffs.
-We establish existence under standard conditions, and the refinement is illustrated in a hegemon–client model of foreign politics.
+We study dynamic relationships in which one party extracts current surplus in ways that degrade the future state, while the counterparty cannot exit but adjusts effort in response.
+Standard stationary Markov equilibria may sustain collapse paths in which short-run extraction dominates strictly positive continuation gains.
+We introduce the Sustainable Exploitation Equilibrium (SEE), a refinement for dynamic games with irreversible failure modeled as an absorbing boundary that eliminates continuation value.
+When a survival-preserving action exists and failure destroys future surplus, equilibria assigning positive probability to collapse are sequentially irrational.
+Equilibrium analysis can therefore be restricted, without loss, to continuation-preserving stationary Markov equilibria.
+Within this restricted domain, viable renegotiation-proofness becomes structural: because failure truncates future surplus, any jointly improving survival-preserving deviation is credible prior to collapse.
+SEE selects the viable, renegotiation-stable equilibrium that maximizes the exploiter’s value.
+Existence is established under standard conditions, and the refinement is illustrated in a hegemon–client setting.
 Keywords: Markov perfect equilibrium, equilibrium refinements, viability, renegotiation-proofness, sustainability
-JEL: C72, C73, D74.
+JEL: C72, C73, D74, D90.
 1 Introduction
-Exploitation arises when one agent extracts value from another whose continued functioning is essential for the relationship itself.
-In such dynamic environments, standard equilibrium concepts such as Markov Perfect Equilibrium (MPE) may admit outcomes in which the exploiter optimally extracts too aggressively, driving the state into collapse despite positive continuation payoffs.
-These equilibria rely on the exploiter being indifferent to catastrophic breakdown once it occurs.
-To address this, this paper introduces the Sustainable Exploitation Equilibrium (SEE), a refinement of stationary Markov Perfect Equilibrium that rules out such collapse outcomes by deriving viability from rationality.
-When exiting a sustainability region entails sufficiently large losses for the exploiter—and when a feasible safe action exists that preserves the relationship—any equilibrium placing positive probability on collapse is sequentially irrational.
-Under these conditions, restricting attention to viable stationary equilibria is without loss.
-SEE then further refines the set of rationally viable stationary MPE by imposing renegotiation-proofness and exploiter-optimal selection.
-SEE is motivated by settings where one party cannot exit but remains responsive, such as in hegemon–client dynamics, as the exploitee retains a strategic effort margin that shapes both payoffs and dynamics.
-In this way, SEE formalizes the idea of “sustainable exploitation”: the exploiter extracts as much as possible while ensuring the long-run survival and productivity of the exploitee.
-This rational-viability foundation distinguishes SEE from equilibrium selection rules that impose sustainability as an external constraint.
+Exploitation arises when one party extracts value from another whose continued functioning is necessary for the relationship itself.
+In many economic and political environments, aggressive extraction can erode the very state that sustains the interaction—whether by depleting productive capacity, destabilizing institutions, or undermining cooperation.
+A central strategic tension therefore arises: the actor with extraction power benefits from pushing the system harder today, but doing so risks destroying the future surplus the relationship generates.
+Standard equilibrium concepts such as Markov Perfect Equilibrium (MPE) do not fully resolve this tension.
+Because equilibria are evaluated over all admissible state transitions, they may sustain collapse paths in which short-run extraction dominates strictly positive continuation values.
+This paper introduces the Sustainable Exploitation Equilibrium (SEE), a refinement of stationary MPE for environments with irreversible failure.
+Why the refinement is necessary under irreversible failure.
+In dynamic environments without absorbing collapse, continuation payoffs are defined on a common domain across strategies.
+Irreversible failure changes this structure: once the state exits the survival region, the continuation problem collapses.
+Standard equilibrium notions evaluate strategies without restricting the admissible continuation domain, thereby permitting profiles that rely on the destruction of future surplus along the equilibrium path.
+SEE makes this restriction explicit by requiring that equilibrium paths preserve the continuation domain before selection among them.
+The refinement therefore reflects a consistency requirement imposed by irreversible failure rather than an additional constraint.
 2 Related Literature
@@
 SEE builds on the refinement tradition in dynamic games.
-(MaskinTirole2001a) develop the theory of Markov perfect equilibrium; (Bernheim1987) introduce coalition-proof Nash equilibrium; (Selten1975) trembling-hand perfection; and (FarrellMaskin1989) renegotiation-proofness in repeated games.
+(Maskin and Tirole, 2001) develop the theory of Markov perfect equilibrium; (Bernheim et al., 1987) introduce coalition-proof Nash equilibrium; (Selten, 1975) trembling-hand perfection; and (Farrell and Maskin, 1989) renegotiation-proofness in repeated games.
 Viability and dynamic control.
-The viability constraint draws from (Aubin1991) Viability Theory and constrained Markov decision processes (Altman1999), ensuring state trajectories remain within sustainability sets.
+The viability constraint draws from (Aubin, 1991) Viability Theory and constrained Markov decision processes (Altman, 1999), ensuring state trajectories remain within sustainability sets.
 This connects to the resource economics literature, e.g.
-(Dasgupta1979), (Chichilnisky1997), and (Dockner2000), where long-run viability is central to dynamic exploitation of resources.
+(Dasgupta, 1979), (Chichilnisky, 1997), and (Dockner et al., 2000), where long-run viability is central to dynamic exploitation of resources.
 Political economy and relational contracts.
-SEE also relates to political economy models of extraction and state capacity ((AcemogluRobinson2006); (BesleyPersson2011)) and to relational contracting without full commitment ((ThomasWorrall1988); (Levin2003)).
+SEE also relates to political economy models of extraction and state capacity ((Acemoglu and Robinson, 2006); (Besley and Persson, 2011)) and to relational contracting without full commitment ((Thomas and Worrall, 1988); (Levin, 2003)).
 In these contexts, weaker parties cannot exit but retain effort margins that shape dynamics.
-SEE formalizes the logic of “restrained exploitation” within a Markovian framework.
+SEE formalizes the logic of restrained exploitation within a Markovian framework.
+Comparison.
+In contrast to the above, this paper studies environments in which exit from a viability region constitutes irreversible failure that eliminates continuation value.
+Standard refinements select among equilibria defined on a fixed continuation domain.
+SEE instead restricts the admissible domain by requiring equilibrium paths to preserve the survival region before equilibrium selection.
+Viability and renegotiation-proofness therefore follow from sequential rationality under irreversible failure.
 3 Model and Definition
@@
 Rational interpretation of viability.
-We interpret transitions outside a sustainability set $V\subset S$ as catastrophic breakdown for the exploiter, in the sense that sufficiently small reductions in extraction can strictly improve continuation payoffs by preserving the relationship.
-Under this interpretation, equilibrium behavior that places positive probability on exiting $V$ is incompatible with sequential rationality when collapse entails sufficiently large losses.
+We interpret transitions outside a sustainability set $V\subset S$ as irreversible failure: absorption into a collapse region that eliminates continuation value.
+Under this interpretation, equilibrium behavior that places positive probability on exiting $V$ is incompatible with sequential rationality when failure entails sufficiently large losses and a survival-preserving action exists.
 Definition 3 (Sustainable Exploitation Equilibrium (SEE)).
@@
   2. (ii)
-    Renegotiation-proofness: At any state $s\in V$, there is no alternative viable profile $\tilde{\sigma}\in\mathcal{E}_{V}$ such that:
+    Renegotiation-proofness: At any state $s\in V$, there is no alternative viable profile $\tilde{\sigma}\in\mathcal{E}_{V}$ such that
     $W_{E}(\tilde{\sigma};s)\geq W_{E}(\sigma^{*};s)\quad\text{and}\quad W_{X}(\tilde{\sigma};s)\geq W_{X}(\sigma^{*};s),$
     with at least one inequality strict.
+Renegotiation is modeled in reduced form: because irreversible failure eliminates continuation value, any jointly improving survival-preserving deviation is credible prior to collapse.
+Renegotiation-proofness therefore acts as a stability requirement on continuation paths within the survival domain.
   3. (iii)
-    Exploiter-optimality: Among all renegotiation-proof elements of $\mathcal{E}_{V}$, $\sigma^{*}$ maximizes the exploiter’s value $W_{X}(\cdot;s)$ at the state $s$ (with ties broken lexicographically).
+    Exploiter-optimality: Among all renegotiation-proof elements of $\mathcal{E}_{V}$, $\sigma^{*}$ maximizes the exploiter’s value $W_{X}(\cdot;s)$ at state $s$ (ties broken lexicographically).
 Any profile satisfying (i)–(iii) is a Sustainable Exploitation Equilibrium.
+Multiplicity and uniqueness.
+The viability restriction removes collapse-inducing equilibria, while renegotiation-proofness eliminates survival-preserving equilibria that leave joint surplus unexploited.
+SEE therefore reduces the equilibrium correspondence relative to MPE and may yield uniqueness, although multiplicity can persist when several survival-preserving equilibria generate identical joint values.
 4 Relation to Nash and Markov Perfect Equilibrium
@@
 Sustainable Exploitation Equilibrium (SEE).
-SEE is defined as a viable, renegotiation-proof MPE with exploiter-optimal selection.
-Formally, a stationary Markov profile $(\sigma_{X},\sigma_{E})$ is an SEE if the aforementioned conditions of Definition 3 are satisfied.
-It follows immediately that
+SEE is defined as a viable, renegotiation-proof MPE with exploiter-optimal selection. It follows immediately that
 $\text{SEE}\subseteq\text{MPE}\subseteq\text{SPE}\subseteq\text{NE}.$
+SEE is strictly stronger than renegotiation-proof MPE.
+Renegotiation-proofness selects among equilibria within a given continuation domain but does not restrict that domain itself.
+With an absorbing collapse state, equilibria may drive the system to failure when short-run gains exceed continuation values and yet remain renegotiation-proof because collapse eliminates future surplus.
+SEE excludes such outcomes by restricting attention to survival-preserving paths, thereby refining both equilibrium selection and the admissible continuation domain.
+Similar observations apply to coalition-proof refinements, which eliminate profitable coalitional deviations but likewise evaluate equilibria on a fixed continuation domain.
 5 Existence under rational viability
-We derive viability from rationality by interpreting exits from a sustainability set as catastrophic for the exploiter.
-We interpret the catastrophe penalty $M$ as a reduced-form representation of franchise loss, regime collapse, or irreversible destruction of the exploitee following breakdown of the relationship.
-Assumption (Catastrophic collapse and safe action).
+We derive viability from rationality by interpreting exits from a sustainability set as irreversible failure.
+Economic interpretation of irreversible failure.
+Irreversible failure is modeled as an absorbing boundary that eliminates continuation value.
+Economically this corresponds to events such as sovereign default, political collapse, financial liquidation, or irreversible depletion of productive capacity.
+When the discounted continuation loss exceeds the maximal short-run gain from over-extraction, survival-preserving behavior follows from sequential rationality.
+Assumption (Irreversible failure and safe action).
 Payoffs are bounded: $|u^{X}(s,x,e)|\leq\bar{u}<\infty$.
@@
 $Q\!\left(V\mid s,\underline{x}(s),\sigma_{E}(s,\underline{x}(s))\right)=1.$
+Assumption (Uniform exit risk for non-survival-preserving actions). For each $s\in V$, let
+$\mathcal{X}_{s}^{\,N}\;=\;\{x\in X(s):Q(V\mid s,x,\sigma_{E}(s,x))<1\}$
+denote the set of actions that are not survival-preserving (given the exploitee’s best response). Assume that the minimal probability of exiting $V$ over all such actions is bounded away from zero:
+$p^{\star}\;=\;\inf_{s\in V}\inf_{x\in\mathcal{X}_{s}^{\,N}}\big(1-Q(V\mid s,x,\sigma_{E}(s,x))\big)\;>\;0.$
 Theorem 1 (Existence of SEE under rational viability).
 Suppose the state space $S$ and action sets $X(s),E(s)$ are compact, payoffs are bounded and continuous, and the transition kernel $Q$ satisfies the Feller property.
-Interpret $S\setminus V$ as a collapse region generating a catastrophic loss for the exploiter.
-Then there exists $M^{\star}<\infty$ such that, for all catastrophe penalties $M\geq M^{\star}$, every stationary Markov Perfect Equilibrium of the penalized game
+Interpret $S\setminus V$ as an absorbing failure region that eliminates continuation value.
+Then there exists $M^{\star}<\infty$ such that, for all failure penalties $M\geq M^{\star}$, every stationary Markov Perfect Equilibrium of the penalized game
 $\hat{u}^{X}(s,x,e)\equiv u^{X}(s,x,e)-M\cdot\mathbf{1}\{s^{\prime}\notin V\},\qquad s^{\prime}\sim Q(\cdot\mid s,x,e),$
@@
 Hence a Sustainable Exploitation Equilibrium exists.
-Sketch.
+Proof.
 Under compactness and continuity, stationary Markov Perfect Equilibria exist.
 Fix any $s\in V$.
-If an equilibrium action induces $Q(S\setminus V\mid s,x,e)>0$, the exploiter’s expected one-period payoff is at most $\bar{u}-Mp$ for some $p>0$.
+If an equilibrium action induces $Q(S\setminus V\mid s,x,e)>0$, the exploiter’s expected one-period payoff is at most $\bar{u}-Mp$ for some $p^{*}>0$.
 Under the safe action $\underline{x}(s)$, the expected penalty is zero and the payoff is at least $\bar{u}$.
@@
 ∎
+Finite collapse threshold.
+Because payoffs are bounded and safe actions exist at every state in the survival domain, there exists a finite failure threshold $M^{*}$ such that any action inducing a positive probability of exit is strictly dominated once $M\geq M^{*}$.
+The threshold depends on (i) the maximal one-period gain from over-extraction, (ii) the minimal probability of exit under non-safe actions, and (iii) the discount factor.
+In particular, higher exit risk under over-extraction lowers the required threshold, while greater patience reduces the attractiveness of failure paths.
 6 Example: Foreign Politics (Hegemon–Client SEE)
@@
 $\max_{e\in E(s)}\;\Big\{b(s)-\phi(e)-d(x)+\delta W^{C}\!\big(f(s,e)-h(x)\big)\Big\}.$
-An interior first-order condition is
-$\phi^{\prime}(e^{*})\;+\;\delta\,W^{C}_{s}\!\big(s^{\prime}\big)\,f_{e}(s,e^{*})\;=\;0,\qquad s^{\prime}\equiv f(s,e^{*})-h(x).$
-Under standard concavity/regularity, the implicit function theorem yields $\partial e^{*}/\partial x$.
+Remark. If $W_{C}$ is differentiable, the interior best response satisfies
+$\phi^{\prime}(e^{*})+\delta W_{C}^{\prime}(s^{\prime})f_{e}(s,e^{*})=0.$
 6.5 Leader’s SEE problem
@@
 $V^{H}(s)\;=\;\max_{x\in X(s)}\;\Big\{\pi(x)-k(s,e^{*}(s,x))+\delta\,V^{H}\!\big(f(s,e^{*}(s,x))-h(x)\big)\Big\}$
-subject to viability $x\leq f(s,e^{*}(s,x))-s_{\min}$. Let $\mu\geq 0$ be the multiplier on viability. The Lagrangian is
-$\mathcal{L}(x;s,\mu)\;=\;\pi(x)-k(s,e^{*})+\delta V^{H}(s^{\prime})\;+\;\mu\big(f(s,e^{*})-h(x)-s_{\min}\big),$
-where $e^{*}\equiv e^{*}(s,x)$ and $s^{\prime}\equiv f(s,e^{*})-h(x)$.
-The (interior) first-order condition with envelope terms from $e^{*}(s,x)$ is
-$\pi^{\prime}(x)\;-\;\delta V^{H}_{s}(s^{\prime})\,\big(h^{\prime}(x)-f_{e}(s,e^{*})\,\partial e^{*}/\partial x\big)\;-\;\mu\,\big(h^{\prime}(x)-f_{e}(s,e^{*})\,\partial e^{*}/\partial x\big)\;=\;0,$
-with complementary slackness $\mu\big(f(s,e^{*})-h(x)-s_{\min}\big)=0$.
-Interpretation.
-The hegemon equates marginal extraction benefits $\pi^{\prime}(x)$ to the discounted shadow cost of lost future capacity $h^{\prime}(x)$ net of the induced effort response $f_{e}\partial e^{*}/\partial x$.
-When viability binds, the shadow price $\mu>0$ further restrains extraction.
-6.6 Stationary SEE and regimes
-A stationary SEE consists of $(x^{*}(s),e^{*}(s,x))$ and an invariant distribution supported on $V=\{s\geq s_{\min}\}$. Under rational viability, two canonical regimes arise:
-  • Interior (viability slack, $\mu=0$).
-If the unconstrained fixed point $s^{*}$ of $s^{\prime}=f(s,e^{*}(s,x^{*}))-h(x^{*})$ satisfies $s^{*}>s_{\min}$, the FOCs imply a generalized “golden rule” where effective regeneration $f_{s}+f_{e}\,\partial e^{*}/\partial s$ balances discounting (details depend on primitives).
-  • Boundary (viability binds, $\mu>0$).
-If the unconstrained $s^{*}$ would fall below $s_{\min}$, SEE selects $s^{*}=s_{\min}$ and $x^{*}=f(s_{\min},e^{*}(s_{\min},x^{*}))-s_{\min}$, with $\mu$ set to satisfy the FOC.
-Additional Remark on Outside Option
-One might alternatively assume that the exploitee $E$ has an outside option $U^{out}_{E}(s)$, in which case viability must be supplemented by an individual rationality (IR) constraint: $W^{E}(s)\geq U^{out}_{E}(s)$ for all $s\in V$.
-This yields an even tighter refinement, as it restricts feasible equilibria to those in which the exploitee weakly prefers participation over exit.
+subject to viability $x\leq f(s,e^{*}(s,x))-s_{\min}$. Let $\mu\geq 0$ be the multiplier on viability.
+Additional Implementative Remarks
+Robustness to small outside options.
+The no-exit assumption sharpens the exploitation structure but is not essential.
+Suppose the exploitee has an outside option yielding value at most $\varepsilon>0$ relative to continuation within the survival domain.
+For sufficiently small $\varepsilon$, survival-preserving stationary Markov equilibria and the renegotiation-stable selection vary continuously.
+When collapse losses for the exploiter exceed the maximal short-run gain from over-extraction, the viability restriction continues to bind and SEE converges to the baseline no-exit refinement as $\varepsilon\to 0$.
+Timing and simultaneity.
+The baseline adopts sequential timing in which extraction precedes effort adjustment, but the refinement does not depend on this structure.
+Under simultaneous moves, stationary Markov equilibria remain well-defined, and the survival-domain restriction continues to exclude collapse-inducing profiles.
+The mechanism therefore reflects the truncation of continuation value at failure rather than the timing protocol.
+Empirical operationalization.
+The state $s_{t}$ can be proxied by observable fiscal or administrative capacity indices (e.g.
+World Bank Governance Indicators), extraction $x_{t}$ by documented concession or compliance flows, and $s_{\min}$ by sovereign default or regime-transition events observable in panel data.
+Structural parameters governing persistence and responsiveness can be calibrated using standard approaches in state-capacity models (e.g.
+Besley and Persson (2011)).
+The viability threshold is disciplined by the present value of foregone extraction rents following collapse.
 7 Conclusion
-With a non-exiting but strategically responsive exploitee, SEE refines Markov Perfect Equilibrium by deriving viability from rationality and imposing renegotiation-proof exploiter-optimal selection.
-Unlike MPE, SEE eliminates equilibria in which the exploiter optimally drives the state to collapse despite positive continuation payoffs.
-In equilibrium, SEE yields either an interior “golden-rule”–type extraction path or, when unconstrained extraction would cause collapse, a boundary outcome in which the state is pinned at the viability threshold.
-The framework applies broadly to settings in which one party cannot exit but retains an effort margin—such as colonial relationships, labor arrangements, and locked-in suppliers—providing a compact refinement for analyzing persistent exploitative interactions in economics and political economy.
+Dynamic relationships often involve asymmetric control over current surplus and future viability.
+When extraction degrades the state sustaining the interaction, standard equilibrium concepts may permit collapse paths even when continuation values remain positive.
+We introduce the Sustainable Exploitation Equilibrium (SEE), a refinement of stationary Markov Perfect Equilibrium for environments with irreversible failure.
+Because failure eliminates continuation value, survival-preserving deviations are credible prior to collapse, making renegotiation-proofness endogenous and restricting equilibrium analysis to viable stationary equilibria.
+If unconstrained extraction preserves viability, SEE yields an interior outcome; if it would induce failure, equilibrium instead converges to the viability boundary—the maximal level of exploitation consistent with survival.
+SEE thus provides a compact framework for analyzing persistent exploitative relationships while clarifying how irreversible failure disciplines incentives without eliminating pressures toward maximal sustainable exploitation.